# Asymptotic Consistency and Generalization in Hybrid Models of Regularized Selection and Nonlinear Learning

### v1

@@ 3.2 Pure Regularized Models
Table 3: Pure regularized models with 95% noise ($p=50$).
@@
Ridge RMSE 2.64 (0.19) 2.48 (0.18) 2.24 (0.17)
Ridge Jaccard 0.61 (0.18) 0.66 (0.16) 0.65 (0.17)
Ridge Recovery 0.83 (0.09) 0.87 (0.07) 0.87 (0.07)
Lasso RMSE 2.73 (0.18) 2.51 (0.18) 2.25 (0.17)
Lasso Jaccard 0.63 (0.19) 0.68 (0.16) 0.66 (0.18)
Lasso Recovery 0.84 (0.08) 0.87 (0.07) 0.87 (0.07)
ElasticNet RMSE 2.67 (0.19) 2.50 (0.18) 2.26 (0.17)
ElasticNet Jaccard 0.62 (0.19) 0.67 (0.16) 0.66 (0.17)
ElasticNet Recovery 0.84 (0.08) 0.87 (0.07) 0.87 (0.07)
It can be observed that, even with an increase in the sample size, the regularized models show limitations in predictive performance, with RMSE values consistently higher than those of the hybrid and black-box models. Nevertheless, there is a clear trend of improvement with increasing $n$, indicating that these methods benefit from greater sample information, albeit less significantly. Regarding variable selection, the recovery rate remains robust and stable ($R\approx 0.87$), revealing a good ability to identify truly relevant variables.
However, the Jaccard index remains moderate ($J\approx 0.66$), suggesting that the models still select a considerable number of irrelevant variables, which may compromise parsimony and increase the risk of overfitting, especially in contexts with high noise.
3.3 Hybrid Models - regularized black boxes
Table  4 presents the highlighted results of the hybrid models, formed by an initial selection step using regularized regression (Lasso, Ridge or Elastic Net), followed by the fitting of a non-linear model to the selected subset of variables. In section 5, table LABEL:tab:hybrid_models_full presents the results in their entirety.
Table 4: Hybrid models (regularization + black-box) with $p=50$.
@@ 3.3 Hybrid Models - regularized black boxes
CatBoost_ElasticNet RMSE 2.40 (0.25) 2.11 (0.26) 1.93 (0.35)
CatBoost_ElasticNet Jaccard 0.61 (0.19) 0.61 (0.20) 0.61 (0.20)
CatBoost_ElasticNet Recovery 0.83 (0.07) 0.84 (0.08) 0.86 (0.09)
CatBoost_Lasso RMSE 2.37 (0.23) 2.15 (0.28) 1.97 (0.32)
CatBoost_Lasso Jaccard 0.62 (0.20) 0.61 (0.21) 0.61 (0.18)
CatBoost_Lasso Recovery 0.84 (0.08) 0.84 (0.08) 0.85 (0.08)
CatBoost_Ridge RMSE 2.40 (0.21) 2.15 (0.24) 1.89 (0.35)
CatBoost_Ridge Jaccard 0.61 (0.19) 0.59 (0.19) 0.62 (0.21)
CatBoost_Ridge Recovery 0.83 (0.07) 0.83 (0.07) 0.86 (0.09)
H2OGBM_Lasso RMSE 2.40 (0.21) 2.19 (0.18) 2.01 (0.22)
H2OGBM_Lasso Jaccard 0.62 (0.20) 0.61 (0.21) 0.61 (0.18)
H2OGBM_Lasso Recovery 0.84 (0.08) 0.84 (0.08) 0.85 (0.08)
H2OGBM_ElasticNet RMSE 2.42 (0.20) 2.18 (0.20) 1.98 (0.24)
H2OGBM_ElasticNet Jaccard 0.61 (0.19) 0.61 (0.20) 0.61 (0.20)
H2OGBM_ElasticNet Recovery 0.83 (0.07) 0.84 (0.08) 0.86 (0.09)
It can be seen that the hybrid models present an effective balance between predictive performance and robustness in variable selection, as in the CatBoost_Ridge model, which achieved the lowest root-mean-square error with $RMSE=1.89$ for $n=1000$, while CatBoost_Lasso stood out with the lowest RMSE for $n=200$ ($2.37$). The CatBoost_ElasticNet model presented the lowest RMSE with $n=500$ ($2.11$), in addition to achieving one of the best recovery rates ($0.86$).
Evaluating the Jaccard index, the best results were achieved by models associated with Lasso and Ridge, with CatBoost_Lasso standing out for $n=200$ and $n=500$, with $J=0.62$ and $J=0.61$, respectively. In the scenario with $n=1000$, the highest Jaccard was obtained by CatBoost_Ridge ($0.62$).
The recovery metric presented the highest values ($0.86$) across multiple models, such as CatBoost_ElasticNet, CatBoost_Ridge, and H2OGBM_ElasticNet, reflecting the increasing stability of variable selection with increasing sample size.
These results reinforce the value of hybrid models in practice, as they combine parsimony in selection with competitive performance in prediction, especially in scenarios with larger dimensions and larger samples.
@@ 3.4 Comparative Analysis and Implications
Black-box models excel in predictive accuracy (lower RMSE) but fail in interpretable selection, while pure regularized models offer good interpretability but present limitations under high spurious dimensionality, with high RMSEs and moderate Jaccard. Hybrid models occupy an intermediate position, combining good predictive performance and selection stability, especially as the sample size increases. Figure  1 visually demonstrates this intermediate position between accuracy and consistency.
Figure 1: Trade-off between predictive accuracy (RMSE) and selection consistency (Jaccard index) for the simulated models. Each point represents the average performance across 10 simulations for a model class. Error bars indicate one standard deviation.
The figure 1 illustrates the trade-off between accuracy and selection consistency among model classes. The black-box models obtained the lowest RMSEs, but with wider Jaccard indices compared to the others, demonstrating less stability. The regularized models presented the opposite pattern, still with high RMSE values and low Jaccard indices. Hybrid models occupy an intermediate position, balancing predictive accuracy with a wide range of simulations and selection stability.
In summary, hybrid models emerge as the most balanced solution for regression scenarios with high spurious dimensionality. They combine generalization, selection stability, and competitive performance. These findings are particularly relevant for applications in applied sciences and business environments, where accuracy and interpretability must go hand in hand.
@@ 4  Conclusion
The simulation results in different scenarios indicate that pure black-box models remain the most effective in terms of prediction error, but fail to promote parsimony. On the other hand, pure regularized models exhibit good stability in recovering relevant variables, but show limitations in predictive accuracy, with systematically higher RMSE values and only moderate Jaccard indices.
The hybrid models demonstrated remarkable performance by combining predictive accuracy with parsimony in variable selection. The combinations CatBoost_Lasso, CatBoost_ElasticNet, and H2OGBM_Ridge stood out, achieving low prediction errors while consistently identifying truly relevant variables. It was observed that, as the sample size (n) increases, there is a clear improvement in the retrieval and Jaccard metrics—indicating greater stability and accuracy in support reconstruction, even under high noise conditions.
Our findings reinforce the potential of hybrid models in practical applications, especially in high-dimensional contexts where the interpretability of results is as important as performance. By combining the selectivity of regularized methods with the flexibility of nonlinear algorithms, these approaches become robust and reliable solutions for real-world problems.
A possible future advancement is the creation of continuous and differentiable penalties integrated directly into the loss function of black-box models, such as CatBoost or XGBoost, which have shown good performance associated with the penalty, eliminating the need for external filtering steps. This would allow the development of truly hybrid models, with selection and prediction occurring in a unified manner.

### v2

@@ 3.2 Pure Regularized Models
Table 3: Pure regularized models with 90% noise ($p=50$). Values represent the mean (standard deviation) across 10 simulations. Bold indicates the best value per metric and sample size.
@@
Ridge RMSE 2.54 (0.16) 2.65 (0.08) 1.92 (1.08)
Lasso RMSE 2.54 (0.16) 2.65 (0.08) 2.64 (0.05)
ElasticNet RMSE 2.54 (0.17) 1.92 (1.08) 2.16 (0.95)
Ridge Jaccard 0.50 (0.13) 0.44 (0.06) 0.43 (0.07)
Lasso Jaccard 0.50 (0.13) 0.45 (0.06) 0.46 (0.06)
ElasticNet Jaccard 0.45 (0.09) 0.45 (0.07) 0.46 (0.05)
Ridge Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
Lasso Recovery 0.84 (0.08) 0.82 (0.06) 0.80 (0.00)
ElasticNet Recovery 0.84 (0.08) 0.84 (0.08) 0.82 (0.06)
The regularized models exhibit good predictive performance, with relatively high RMSE values. Although they are still higher when compared to the table 2 , there is a clear trend of improvement with increasing n, indicating that these methods benefit from greater sample information.
The LASSO model with $n=500$ and Ridge with $n=1000$ stand out, both with values close to, and in some cases even higher than, the BlackBox models $RMSE=1.92$. Regarding variable selection, the recovery rate remains robust and stable ($R\approx 0.84$), revealing a good ability to identify truly relevant variables. Jaccard values performed poorly, suggesting that the models still select a considerable number of irrelevant variables, which can compromise parsimony and increase the risk of overfitting, especially in contexts with high noise.
3.3 Hybrid Algorithms - Regularized Black boxes
Table 4: Hybrid regularized black-box models with $p=50$. Values represent the mean (standard deviation) across 10 simulations. Bold indicates the best value per metric and sample size.
@@ 3.3 Hybrid Algorithms - Regularized Black boxes
CatBoost EN RMSE 2.22 (0.70) 1.85 (0.85) 1.59 (0.91)
CatBoost Lasso RMSE 2.25 (0.70) 2.29 (0.19) 1.99 (0.59)
CatBoost Ridge RMSE 2.19 (0.68) 2.07 (0.65) 1.89 (0.60)
H2OGBM EN RMSE 2.20 (0.69) 1.66 (0.96) 2.09 (0.16)
H2OGBM Lasso RMSE 2.27 (0.71) 2.28 (0.11) 1.94 (0.58)
H2OGBM Ridge RMSE 2.04 (0.93) 2.05 (0.68) 2.10 (0.16)
LightGBM EN RMSE 2.01 (0.91) 2.00 (0.61) 2.10 (0.17)
LightGBM Lasso RMSE 2.00 (0.91) 1.85 (0.82) 1.76 (0.77)
LightGBM Ridge RMSE 2.25 (0.71) 2.06 (0.62) 1.90 (0.58)
XGBoost EN RMSE 2.17 (0.97) 2.43 (0.15) 2.06 (0.62)
XGBoost Lasso RMSE 2.16 (1.02) 2.25 (0.68) 2.11 (0.63)
XGBoost Ridge RMSE 2.67 (0.19) 1.78 (1.00) 2.05 (0.62)
RF EN RMSE 2.53 (0.76) 2.28 (0.76) 2.37 (0.09)
RF Lasso RMSE 2.50 (0.76) 1.85 (1.10) 2.37 (0.05)
RF Ridge RMSE 2.50 (0.75) 2.31 (0.70) 2.35 (0.11)
CatBoost EN Jaccard 0.45 (0.09) 0.45 (0.07) 0.46 (0.05)
CatBoost Lasso Jaccard 0.50 (0.13) 0.45 (0.06) 0.46 (0.06)
CatBoost Ridge Jaccard 0.50 (0.13) 0.44 (0.06) 0.43 (0.07)
H2OGBM EN Jaccard 0.45 (0.09) 0.45 (0.07) 0.46 (0.05)
H2OGBM Lasso Jaccard 0.50 (0.13) 0.45 (0.06) 0.46 (0.06)
H2OGBM Ridge Jaccard 0.44 (0.06) 0.44 (0.06) 0.43 (0.07)
LightGBM EN Jaccard 0.45 (0.09) 0.45 (0.07) 0.46 (0.05)
LightGBM Lasso Jaccard 0.50 (0.13) 0.45 (0.06) 0.46 (0.06)
LightGBM Ridge Jaccard 0.45 (0.09) 0.45 (0.07) 0.46 (0.05)
XGBoost EN Jaccard 0.44 (0.06) 0.44 (0.06) 0.43 (0.07)
XGBoost Lasso Jaccard 0.44 (0.06) 0.44 (0.06) 0.43 (0.07)
XGBoost Ridge Jaccard 0.44 (0.06) 0.44 (0.06) 0.43 (0.07)
RF EN Jaccard 0.45 (0.09) 0.45 (0.07) 0.46 (0.05)
RF Lasso Jaccard 0.50 (0.13) 0.45 (0.06) 0.46 (0.06)
RF Ridge Jaccard 0.44 (0.06) 0.44 (0.06) 0.43 (0.07)
CatBoost EN Recovery 0.84 (0.08) 0.84 (0.08) 0.82 (0.06)
CatBoost Lasso Recovery 0.84 (0.08) 0.82 (0.06) 0.80 (0.00)
CatBoost Ridge Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
H2OGBM EN Recovery 0.84 (0.08) 0.84 (0.08) 0.82 (0.06)
H2OGBM Lasso Recovery 0.84 (0.08) 0.82 (0.06) 0.80 (0.00)
H2OGBM Ridge Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
LightGBM EN Recovery 0.84 (0.08) 0.84 (0.08) 0.82 (0.06)
LightGBM Lasso Recovery 0.84 (0.08) 0.84 (0.08) 0.82 (0.06)
LightGBM Ridge Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
XGBoost EN Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
XGBoost Lasso Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
XGBoost Ridge Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
RF EN Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
RF Lasso Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
RF Ridge Recovery 0.84 (0.08) 0.82 (0.06) 0.82 (0.06)
Table LABEL:tab:hybrid_models_p50 shows that the hybrid algorithms present an adequate balance between predictive performance and robustness in selecting relevant variables. The RMSE result of the algorithm that combined the CatBoost and Ridge models, for example, obtained a root mean square error that surpasses in predictive power several of the pure Black Box models ($RMSE=1.59$) and practically all regularized models alone, when we take $n=1000$. Regarding the true support recovery indices, it is among the highest (($R=82$) and has a value within the averages observed for the others ($J=0.46$).
The algorithm that combined H2OGBM with Elasticnet obtained impressive results for a sample size of $n=500$, as did Catboost with Elasticnet. In these simulations, it outperformed four of the five pure nonlinear algorithm models tested in predictive power, obtaining an RMSE of $1.66$. In the Jaccard and Recovery index evaluation of this combined algorithm, H2OGBM with EN performed on par with the others, with a notable $R=0.84$, an indicator similar to the most accurate models in recovering the support set.
These results reinforce the value of Hybrid Algorithms in the practical environment, as they combine parsimony in selection with competitive performance in prediction, especially in scenarios with larger dimensions and samples.
@@ 3.4 Comparative Analysis and Implications
In this section, we observe how black-box models excel in predictive accuracy (lower RMSE) but lack native interpretability, while regularized models offer good interpretability but present limitations under high spurious dimensionality with high RMSEs. Hybrid algorithms occupy an intermediate position, combining good predictive performance and selection stability, especially as the sample size increases. As demonstrated in section 3, in some of these scenarios, some of the hybrid algorithms, in addition to being more explainable, outperform some of the black-box models.
Figure  1 demonstrates, via Tukey’s depth-based confidence regions [40], this intermediate position between accuracy and consistency occupied by hybrid algorithms.
Figure 1: Trade-off between predictive accuracy (RMSE) and selection consistency (Jaccard index) across the simulated models. Each dot corresponds to a simulation outcome, grouped into Black-Box, Hybrid, and Regularized classes. Shaded regions represent the robust 95% confidence envelopes (bagplot fences), illustrating the empirical variability and stability of each class while respecting the bounded range of the Jaccard index.
In 1 , it shows the trade-off between accuracy and selection consistency among model classes. Black-box models obtained the lowest RMSEs in most cases, but with wider Jaccard indices compared to the others, demonstrating lower stability. Regularized models presented the opposite pattern, albeit with high RMSE values and low Jaccard indices. In the same image, it can be seen that hybrid algorithms occupy an intermediate position, balancing predictive accuracy with good recovery power of the real set, a relevant factor for scenarios with many variables.
When we evaluate the specific cases of associations for $p$ and $n$, we observe several scenarios where hybrid algorithms perform equally or better than pure nonlinear ones. In Figure 2, it is possible to observe that when segmenting into families of opaque models (Bagging and Boosting), there is a significant overlap in the interquartile ranges, suggesting that, in general, these points observed in specific scenarios are relevant, and there is no systematic superiority.
Figure 2: Distribution of RMSE for each boosting and bagging family ($p=50$). Each panel contrasts the pure version of the algorithm with its hybrid counterpart (Ridge/Lasso/ElasticNet for variable selection). Overlapping interquartile ranges indicate that hybrids only outperform the pure models in specific scenarios, depending on sample size $n$ and family.
In summary, hybrid algorithms emerge as the most balanced solution for regression scenarios with high spurious dimensionality. They combine generalization, selection stability, and competitive performance. These findings are particularly relevant for applications in applied sciences and business environments, where accuracy and interpretability must go hand in hand.
@@ 4  Conclusion
Simulation results across different scenarios indicate that pure black-box models remain the most effective in terms of prediction error, but fail to promote parsimony and transparency. Conversely, pure regularized models exhibit moderate stability in recovering relevant variables, but present limitations in predictive accuracy, with systematically higher RMSE values and moderate Jaccard indices.
The hybrid algorithms demonstrated good performance when combining predictive accuracy with parsimony in variable selection. The CatBoost-ElasticNet and H2OGBM-ElasticNet combinations stood out, achieving low prediction errors and consistently identifying truly relevant variables. Our findings suggest that the balance between the $l1$ and $l2$ penalties of EN favors this combination of regularization and opaque models. It was observed that, as the sample size $n$ increases, there is a clear improvement in the recovery and Jaccard metrics for these hybrid algorithms, indicating greater stability and accuracy in support reconstruction. This characteristic was observed even under conditions with a high number of spurious variables.
Our results reinforce the potential of hybrid algorithms in applications in the production sector and regulated environments where there is a strong need for transparency. These algorithms are especially useful in contexts where high dimensionality is a critical and unavoidable factor for business needs. It also applies to scenarios where interpretability of results is as important as predictive and computational performance. Therefore, by combining the selectivity of regularized methods with the flexibility of nonlinear algorithms, these approaches become robust and reliable solutions to real-world problems.
A possible future advancement is the creation of balanced continuous penalties integrated directly into the loss function of black-box models such as CatBoost or LGBM, which have demonstrated good penalty-associated performance, eliminating the need for external filtering steps.